OpenStudy (anonymous):

y = 2x + 4 2x - y = 6

OpenStudy (anonymous):

y = 2x + 4 2x - y = 6 . Solve the system of equations using substitution. A) (2,8) B) (2,-8) C) (10,24) D) no solution

OpenStudy (anonymous):

2x-y =6 if we move y to other side and 6 to other side; 2x-6 =y

OpenStudy (anonymous):

2x-6 =y 2x +4 = y ; if we sub the first from the second ; -10 =0; this mean there is not solution to this equations

OpenStudy (anonymous):

answer is D

OpenStudy (anonymous):

thxs